INAF queue observing, 2016 Jan 10-11

Observer: R. Carini, A. Marchetti
Telescope Operator: Steve Allanson
Support Astronomer: D. Thompson on call
Instrument:

Summary

There is snow ice on the roof (some pictures : https://drive.google.com/folderview?id=0B7TYElRMoLdGaEFfRjdESmlOZms&usp=sharing), there is fog with intermittent snow. The forecast has > 0% chance of snow through 4am. Humidity is at 92% currently. The dome is closed tonight for icy enclosure + weather.

Night duration (12\xB0 el twilight) 19:00-6:00 ( 02:00am-01:00pm UT) = 11 h 0m

Observing time: 0 hrs 0 m

Weather time loss: 11 hrs 0m (100%)

technical time loss 0 m

Data Summary

Project_RUN Objects Filter ExpTime Notes

Details

Article text.

-- %USERSIG{RobertaCarini - 2016-01-10}%

Comments

 
Topic revision: r7 - 05 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback